OpenStudy (anonymous):

Solve for x. x^2 - 2x - 24 = 0 Select one of the options below as your answer: A. -4, -6 B. -4, 6 C. 2, -6 D. 4, 6

OpenStudy (anonymous):

(x-6)(x+4)=0 x=6 or -4

OpenStudy (anonymous):

factor it out x^2-2x-24=(x-6)(x+4)=0 x=6,-4

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

it is not A it is B